OpenStudy (anonymous):

Find the quadratic polynomial whose graph goes through the points (-1,4), (0,6), and (1,12). f(x)= x^2+ x+

OpenStudy (anonymous):

y=(x-a)^2+h 4=r(-1-a)^2+h 6=r(0-a)^2+h 12=r(1-a)^2+h

OpenStudy (anonymous):

solve for r,a,and h

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

someone help me with the ans, i keep getting the wrong ones..

OpenStudy (anonymous):

is that last part what the equation will look like? : f(x)= x^2+ x+

OpenStudy (anonymous):

it is not very hard to solve it by hand but it is long and tedious

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@rukh , is that last part, "f(x)= x^2+ x+", what the equation will look like?

OpenStudy (anonymous):

y=r(x-a)^2+h

OpenStudy (anonymous):

substitute the number in for r,a,h in

OpenStudy (anonymous):

ok